OpenStudy (anonymous):

A rectangular garden has an area of 24 square meters. The width is 8 meters less than twice the length. Find the dimensions of the garden.

OpenStudy (shamil98):

l(w) = 24 w = 2L - 8 L(2L - 8) = 24 2L^2 - 8L - 24 = 0 L^2 - 4L - 12 = 0 (L - 6)(L - 2) = 0 L = 6 , L = 2 6w = 24 w = 4 or 2w = 24 w = 12

OpenStudy (shamil98):

Oh, oops i made a factoring mistake it's (L - 6)(L+2) = 0 L = 6 , L = -2 you can't have a negative length.. so its just L = 6 plugging it in 6w = 24 w = 4. L = 6, W = 4